At t = 0 the hydrogen atom is in the superposition state Ξ¨(π‘Ÿβƒ—, 0) =1√3πœ“100 + π΄πœ“210 where A is a real positive constant. Find A by normalization


1
Expert's answer
2020-08-02T15:04:32-0400

Given wave function is

"\psi(r,0) = \frac{1}{\sqrt {3}} \psi _{100}+ A \psi_{210}"

According to Normalization condition,

"\int |\psi(r,0)|^2 dr=1"

"\int |\frac{1}{\sqrt {3}} \psi _{100}+ A \psi_{210}|^2 dr=1"

"\frac{1}{3}\int |\psi _{100}|^2 dr + A^2 \int |\psi _{210}|^2 dr=1"

"\frac{1}{3}+A^2=1"

Since "\int |\psi _{100}|^2 dr =1 and" "\int |\psi _{210}|^2 dr=1"

On simplifying above equation, we get

"A=\sqrt\frac{2}{3}"